### v3

# On schurity of the dihedral group $\boldsymbol{D_{2p}}$

Abstract.
A finite group $G$ is called a Schur group if every $S$-ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of $\operatorname{Sym}(G)$ that contains all right translations. One of the crucial open questions in the $S$-ring theory is the question on existence of an infinite family of nonabelian Schur groups. We prove that a dihedral group of order $2p$, where $p$ is a Fermat prime or prime of the form $p=4q+1$, where $q$ is also prime, is Schur. Towards this result, we prove nonexistence of a difference set in a cyclic group of order $p\neq 13$ and classify all $S$-rings over some dihedral groups.
Keywords: Schur rings, Schur groups, difference sets.
MSC: 05E30, 05B10, 20B25.
The author is supported by The Israel Science Foundation (project No. 87792731)
1. Introduction
A Schur ring or $S$-ring over a finite group $G$ can be defined as a subring of the group ring $\mathbb{Z}G$ that is a free $\mathbb{Z}$-module spanned by a partition of $G$ closed under taking inverse and containing the identity element $e$ of $G$ as a class. The first construction of such ring was proposed by Schur [32]. The general theory of $S$-rings was developed by Wielandt [34]. Schur and Wielandt used $S$-rings to study permutation groups containing regular subgroups. To the present moment, $S$-rings have become a useful tool for studying Cayley graphs, in particular, the Cayley graph isomorphism problem. On the other hand, the theory of $S$-rings continues to develop itself. Concerning the theory of $S$-rings, we refer the readers to the survey [22].
One of the crucial properties of $S$-rings is the schurity property. If $K$ is a subgroup of the symmetric group $\operatorname{Sym}(G)$ containing the group $G_{r}$ of all right translations, then the partition of $G$ into the orbits of the stabilizer $K_{e}$ defines an $S$-ring over $G$. As Wielandt wrote [35], Schur had conjectured that every $S$-ring can be constructed in such way from a suitable permutation group. However, this conjecture was disproved by Wielandt [34]. In honor of this Schur’s fallacy, $S$-rings arising from permutation groups were called schurian, whereas groups whose all $S$-rings are schurian were called Schur [27]. The schurity property for $S$-rings is closely related to the isomorphism poblem for Cayley graphs (see [6, 22]). The problem of determining of all Schur groups was posed by Pöschel [27]. In general, this problem seems hard, in particular, because the number of $S$-rings over a given group is exponential in the order of this group.
The most of the results on Schur groups are concerned with abelian groups. They can be found in [10, 11, 12, 19, 23, 25, 27, 30]. Necessary conditions of schurity for nonabelian groups were obtained in [26]. In [31], it was shown that every nilpotent Schur group belongs to one of the explicitly given infinite families of groups. Nonschurity of some nonabelian groups was checked in [23, 29]. Computer calculations [36] imply that there are nonabelian Schur groups. The largest known nonabelian Schur group has order $63$. It looks natural to ask the following question.
Question.
Does an infinite family of nonabelian Schur groups exist?
The above question seems to be difficult. One reason for this is that the theory of $S$-rings over nonabelian groups is almost undeveloped. Another one is that attempts to find the required family meet some problems which are hard in general, e.g. the problems of existence of a difference set and computation of cyclotomic numbers. The first of them will be discussed further.
In this paper, we are interested in schurity of dihedral groups. Schurity of dihedral $2$-groups was studied in [23]. Namely, in that paper all $S$-rings of rank at most $5$ over dihedral $2$-groups were classified and it was shown that only $S$-rings associated with divisible difference sets can be nonschurian among them. More precisely, in the present paper we are interested in schurity of a dihedral group of order $2p$, where $p$ is prime. In [26], it was proved that such group is not Schur whenever $p\equiv 3\mod~{}4$ and $p>11$. The main result of this paper is the following theorem which is the first noncomputational result stating an existence of nonabelian Schur groups.
Theorem 1.1.
Let $p$ be a prime. If $p$ is a Fermat prime or $p=4q+1$, where $q$ is a prime, then a dihedral group of order $2p$ is Schur.
Towards the proof of Theorem 1.1, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime, i.e. a prime of the form $p=2^{k}+1$, where $k\geq 1$, or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$ (Theorem 4.1). Recall that primes of the form $12q+$ 1, where $q$ is also prime, are known as safe primes (see, e.g., [28]).
It is not known whether there are infinitely many Fermat primes or primes of the form $rq+1$, where $q$ is also prime and $r\in\{2,4\}$. Only five Fermat primes are known; the largest of them is $65537$. The question whether a number of primes of the form $rq+1$ is infinite is closely related to the famous Dickson conjecture [7] (see also [28, p. 372]). Namely, if the Dickson conjecture is true, then there are infinitely many primes of the form $14q+$ 1. It should be mentioned that the Dickson conjecture is widely believed. For example, the Green-Tao theorem [14] stating that the sequence of prime numbers contains arbitrarily long arithmetic progressions can be considered as a great support to the Dickson conjecutre. Note that the discussed family of primes is infinite modulo the Bateman-Horn conjecture (see [1, 17]) or the generalized Hardy-Littlewood conjecture (see [15]). Computational results on primes $q$ such that $12q+$ 1 or $14q+$ 1 is also prime can be found on the web-page [24].
One of the crucial questions which appears during the studying of schurity of dihedral groups is the question on existence of a difference set in a cyclic group (see [23, 26]). A subset $D$ of the group $G$ is called a difference set if there is a nonnegative integer $\lambda$ such that every nonidentity element $g\in G$ has exactly $\lambda$ representations in the form $g=d_{1}^{-1}d_{2}$, $d_{1},d_{2}\in D$. Difference sets were introduced by Singer [33] and systematic studying of difference sets was started with the papers of Hall [16] and Bruck [5]. For details on difference sets, we refer the readers to [3, Chapter VI]. The keynote ingredient in the proof of Theorem 1.1 is the following theorem.
Theorem 1.2.
Let $p$ be a prime. If $p$ is a Fermat prime or $p=4q+1$, where $q>3$ is a prime, then there is no a nontrivial difference set in a cyclic group of order $p$.
Note that Theorem 1.2 for a Fermat prime $p$ is very easy and immediately follows from a simple counting argument, whereas the proof of this theorem for $p=4q+1$ requires some technique of working with multipliers of difference sets. The condition $q>3$ in Theorem 1.2 is essential because there is a nontrivial difference set in a cyclic group of order $13$ (see Remark $1$).
We finish the introduction with a brief outline of the paper. Section $2$ contains a necessary background of $S$-rings. In Section $3$, we provide basic definitions and facts concerned with difference sets and prove Theorem 1.2. In Section $4$, we classify all $S$-rings over a dihedral group of order $2p$, where $p$ is a Fermat prime or $p=rq+1$ for a prime $q$ and $r\in\{2,4\}$, and prove Theorem 1.1. Section $5$ contains some further remarks.
The author would like to thank Prof. Ilia Ponomarenko for a comment on Lemma 4.3 and Prof. Gareth Jones for a comment on the Bateman-Horn conjecture. The author is also grateful to Dr. Saveliy Skresanov for a comment on Lemma 5.1.
2. $S$-rings
In this section, we follow [23] in general. Let $G$ be a finite group and $\mathbb{Z}G$ the integer group ring. The identity element and the set of all non-identity elements of $G$ are denoted by $e$ and $G^{\#}$, respectively. The symmetric group of the set $G$ is denoted by $\operatorname{Sym}(G)$. If $K\leq\operatorname{Sym}(G)$, then the set of all orbits of $K$ on $G$ is denoted by $\operatorname{Orb}(K,G)$. The subgroup of $\operatorname{Sym}(G)$ induced by the right multiplications of $G$ is denoted by $G_{r}$. If $X\subseteq G$, then the element $\sum\limits_{x\in X}{x}$ of the group ring $\mathbb{Z}G$ is denoted by $\underline{X}$. The set $\{x^{-1}:x\in X\}$ is denoted by $X^{-1}$.
A subring $\mathcal{A}\subseteq\mathbb{Z}G$ is called an $S$-ring (a Schur ring) over $G$ if there exists a partition $\mathcal{S}=\mathcal{S}(\mathcal{A})$ of $G$ such that:
$(1)$ $\{e\}\in\mathcal{S}$;
$(2)$ if $X\in\mathcal{S}$, then $X^{-1}\in\mathcal{S}$;
$(3)$ $\mathcal{A}=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:\ X\in\mathcal{S}\}$.
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, $\mathbb{Z}G$ is an $S$-ring. The partition $\{\{e\},G\setminus\{e\}\}$ defines the $S$-ring of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$.
A set $T\subseteq G$ is called an $\mathcal{A}$-set if $\underline{T}\in\mathcal{A}$. If $T$ is an $\mathcal{A}$-set, then put $\mathcal{S}(\mathcal{A})_{T}=\{X\in\mathcal{S}(\mathcal{A}):~{}X\subseteq T\}$. A subgroup $A\leq G$ is called an $\mathcal{A}$-subgroup if $A$ is an $\mathcal{A}$-set. The $S$-ring $\mathcal{A}$ is called primitive if there are no nontrivial proper $\mathcal{A}$-subgroups of $G$ and imprimitive otherwise. One can verify that for every $\mathcal{A}$-set $X$, the groups $\langle X\rangle$ and $\operatorname{rad}(X)=\{g\in G:\ gX=Xg=X\}$ are $\mathcal{A}$-subgroups.
Let $\{e\}\leq A\unlhd B\leq G$. A section $B/A$ is called an $\mathcal{A}$-section if $B$ and $A$ are $\mathcal{A}$-subgroups. If $S=B/A$ is an $\mathcal{A}$-section, then the module
$\mathcal{A}_{S}=Span_{\mathbb{Z}}\left\{\underline{X}^{\pi}:~{}X\in\mathcal{S}(\mathcal{A}),~{}X\subseteq B\right\},$
where $\pi:B\rightarrow B/A$ is the canonical epimorphism, is an $S$-ring over $S$.
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that $\underline{X}\cdot\underline{Y}=\sum\limits_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}$. Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$. The $S$-ring $\mathcal{A}$ is called commutative if $c_{XY}^{Z}=c_{YX}^{Z}$ for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$. If $\mathcal{A}$ is symmetric, then $\mathcal{A}$ is commutative. The following equalities hold
(1) $|Z|c^{Z^{-1}}_{XY}=|X|c^{X^{-1}}_{YZ}=|Y|c^{Y^{-1}}_{ZX},$
(2) $\sum\limits_{Y\in\mathcal{S}(\mathcal{A})}c_{XY}^{Z}=|X|,$
for all $X,Y,Z\in\mathcal{S}(\mathcal{A})$ (see [6, Proposition 2.1.12]).
Let $\mathcal{A}$ and $\mathcal{A}^{\prime}$ be $S$-rings over groups $G$ and $G^{\prime}$, respectively. A bijection $f$ from $G$ to $G^{\prime}$ is called a (combinatorial) isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$ if for every $X\in\mathcal{S}(\mathcal{A})$ there is $X^{\prime}\in\mathcal{S}(\mathcal{A}^{\prime})$ such that $\{(x^{f},y^{f})\in G\times G:~{}yx^{-1}\in X\}=\{(x^{\prime},y^{\prime}):~{}y^{\prime}(x^{\prime})^{-1}\in X^{\prime}\}$. If there is an isomorphism from $\mathcal{A}$ to $\mathcal{A}^{\prime}$, we say that $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic.
A bijection $f\in\operatorname{Sym}(G)$ is defined to be a (combinatorial) automorphism of $\mathcal{A}$ if for all $x,y\in G$, the basic sets containing the elements $yx^{-1}$ and $y^{f}(x^{-1})^{f}$ coincide. The set of all automorphisms of $\mathcal{A}$ forms a group called the automorphism group of $\mathcal{A}$ and denoted by $\operatorname{Aut}(\mathcal{A})$. One can see that $\operatorname{Aut}(\mathcal{A})\geq G_{r}$ and $\mathcal{A}$ is isomorphic to an $S$-ring over a group $H$ if and only if $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $H$. A group automorphism $\varphi\in\operatorname{Aut}(G)$ is defined to be a Cayley automorphism of $\mathcal{A}$ if $X^{\varphi}=X$ for every $X\in\mathcal{S}(\mathcal{A})$. It is easy to check that $\varphi\in\operatorname{Aut}(\mathcal{A})$.
Let $K$ be a subgroup of $\operatorname{Sym}(G)$ containing $G_{r}$. The $\mathbb{Z}$-submodule
$V(K,G)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~{}X\in\operatorname{Orb}(K_{e},~{}G)\},$
where $K_{e}$ is a stabilizer of $e$ in $K$, is an $S$-ring over $G$ as it was proved by Schur [32]. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. If $\operatorname{rk}(\mathcal{A})=2$, then $\mathcal{A}=V(\operatorname{Sym}(G),G)$. So the $S$-ring of rank $2$ is schurian. Clearly, two isomorphic $S$-rings are schurian or not simultaneously.
Let $K\leq\operatorname{Aut}(G)$. Then $\operatorname{Orb}(K,G)$ forms a partition of $G$ that defines an $S$-ring $\mathcal{A}$ over $G$. In this case, $\mathcal{A}$ is called cyclotomic and denoted by $\operatorname{Cyc}(K,G)$. It is easy to see that $\operatorname{Cyc}(K,G)=V(G_{r}K,G)$ and hence every cyclotomic $S$-ring is schurian. The next lemma follows from the results of [27] (see also [25, Lemma 2.4]).
Lemma 2.1.
Every $S$-ring over a group of prime order is cyclotomic and hence schurian.
Let $H$ be a normal $\mathcal{A}$-subgroup of $G$. The $S$-ring $\mathcal{A}$ is called the wreath product of $\mathcal{A}_{H}$ and $\mathcal{A}_{G/H}$ if every basic set $X$ of $\mathcal{A}$ outside $H$ is a union of some $H$-cosets or, equivalently, $H\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus H}$.
Lemma 2.2.
[23, Theorem 3.2] The wreath of two $S$-rings is schurian if and only if each of the operands so is.
3. Difference sets
A subset $D$ of $G$ is called a difference set in $G$ if
$\underline{D}\cdot\underline{D}^{-1}=ke+\lambda\underline{G}^{\#},$
where $k=|D|$ and $\lambda$ is a positive integer. The numbers $(v,k,\lambda)$, where $v=|G|$, are called the parameters of $D$. It is easy to check that $G\setminus D$ is a difference set with parameters $(v,v-k,v-2k+\lambda)$. The empty set, the set $G$, and every subset of $G$ of size $1$ or $|G|-1$ are difference sets. The difference set $D$ is called nontrivial if $22\leq|D|\leq|G|-$ 2. A simple counting argument implies that
(3) $k(k-1)=(v-1)\lambda.$
For the general theory of difference sets, we refer the readers to [2] and [3, Chapter VI].
Lemma 3.1.
Let $p$ be a prime of the form $p=2q+1$, where $q$ is an odd prime, and $D$ a nontrivial difference set in a cyclic group of order $p$. Then $D$ has Paley parameters $(2q+1,q,\frac{q-1}{2})$ or $(2q+1,q+1,\frac{q+1}{2})$.
Proof.
Eq. (3) implies that $k(k-1)=2q\lambda$. Therefore $k$ or $k-1$ is divided by $q$. Since $D$ is nontrivial, we obtain that $k\in\{q,q+1\}$. Now the required follows from Eq. (3). ∎
Note that $p\equiv 3\mod~{}4$ in Lemma 3.1 because $q$ is odd. For each prime $p$ of the form $p=2q+1$, where $q$ is an odd prime, there is a Paley difference set in a cyclic group of order $p$ with parameters $(2q+1,q,\frac{q-1}{2})$ (see, e.g., [3, Section VI.8]).
Let $G$ be abelian. Following [13], we say that an integer $t$ coprime to $v$ is a (numerical) multiplier of $D$ if $D^{(t)}=\{d^{t}:~{}d\in D\}=Dg=\{dg:~{}d\in D\}$ for some $g\in G$. The set of automorphisms $\sigma_{t}:g\rightarrow g^{t}$, $g\in G$, where $t$ runs over all multipliers of $D$, forms a subgroup of $\operatorname{Aut}(G)$ denoted by $M(D)$. For more information on multipliers of difference sets, we refer the readers to [13].
Proof of Theorem 1.2.
Let $A$ be a cyclic group of order $p$. Assume that there is a nontrivial difference set $D$ in $A$. Let $(p,k,\lambda)$ be the parameters of $D$. We may assume that $k<\frac{p}{2}$. Indeed, for otherwise $A\setminus D$ is a difference set with $|A\setminus D|<\frac{p}{2}$ and we replace $D$ by $A\setminus D$. If $p$ is a Fermat prime, then $p-1$ is a $2$-power. Since $k$ and $k-1$ are coprime, Eq. (3) yields that $k=p$ or $k=p-1$, a contradiction to nontriviality of $D$.
Now let $p=4q+1$, where $q>3$ is prime.
Lemma 3.2.
In the above notations, $D$ has parameters $(4q+1,q,\frac{q-1}{4})$ or $(4q+1,q+1,\frac{q+1}{4})$.
Proof.
Eq. (3) implies that $k(k-1)=4q\lambda$. Therefore $k$ or $k-1$ is divided by $q$. Since $D$ is nontrivial and $k<\frac{p}{2}$, we conclude that
$k\in\{q,q+1,2q\}.$
If $k=2q$, then $2q-1=2\lambda$ by Eq. (3). We obtain a contradiction because $12q-$ 1 is odd, whereas $2\lambda$ is even. Thus, $k=q$ or $k=q+1$. Now from Eq. (3) it follows that $\lambda=\frac{q-1}{4}$ in the former case and $\lambda=\frac{q+1}{4}$ in the latter one. ∎
Let $M=M(D)$. Clearly, $M\leq\operatorname{Aut}(A)\cong C_{p-1}=C_{4q}$ and hence $|M|$ divides $4q$. Suppose that $|M|$ is even. Then $M$ contains an automorphism of order $2$, i.e. $\sigma_{0}:a\rightarrow a^{-1}\in\operatorname{Aut}(A)$. However, this contradicts to [2, Theorem 3.3] which states that $-1$ is never a multiplier of a nontrivial difference set in a cyclic group. Therefore, $|M|$ is odd and hence
$|M|\in\{1,q\}.$
Suppose that $|M|=1$, i.e. $M$ is trivial. Let $t$ be a positive integer such that $t<p$ and the automorphism $\sigma_{t}:a\rightarrow a^{t}$ of $A$ is of order $q$. We are going to prove that $t$ is a nontrivial multiplier of $D$ and by that to obtain a contradiction. Let $n=k-\lambda$. From Lemma 3.2 it follows that $n>1$ and $n<p$. The latter yields that $\operatorname{GCD}(p,n)=1$. Let $u$ be a prime divisor of $n$. There exists an automorphism $\sigma_{u}$ of $A$ such that $\sigma_{u}:a\rightarrow a^{u}$ because $u\leq n<p$. If $|\sigma_{u}|$ is even, then the group $\langle\sigma_{u}\rangle$ contains $\sigma_{0}:a\rightarrow a^{-1}$ and hence $u^{i}\equiv-1\mod~{}p$ for some integer $i$. Therefore $u$ is self-conjugate in the sense of [13, p. 7]. Suppose that $|\sigma_{u}|$ is odd. Then $|\sigma_{u}|=q$ because $|\sigma_{u}|$ divides $|\operatorname{Aut}(A)|=p-1=4q$. Together with $|\sigma_{t}|=q$, this implies that $t\equiv u^{i}\mod~{}p$ for some integer $i$. Finally, observe that
$\frac{n}{\operatorname{GCD}(p,n)}=n=k-\lambda>\lambda$
because $\lambda<\frac{k}{2}$. Thus, all conditions of [13, Theorem 3.1] hold for $n_{1}=n$ and $t$ and hence $t$ is a multiplier of $D$.
Now suppose that $|M|=q$. From [3, VI.Theorem 2.6] it follows that there exists a translate of $D$ which is $M$-invariant. Each translate of $D$ is also a difference set in $A$ with the same parameters as $D$ because $A$ is abelian. So without loss of generality, we may assume that $D$ is $M$-invariant. This means that $D$ is a union of some orbits of $M$. Each non-identity orbit of $M$ is a cyclotomic class of order $4$ (see [3, p. 357] for the definition). Recall that $|D|=k\in\{q,q+1\}$ by Lemma 3.2. If $|D|=q$, then $D$ is a non-identity orbit of $M$; if $|D|=q+1$, then $D$ is a union of a non-identity orbit of $M$ with $\{e\}$.
The number $p$ can be presented in the form
(4) $p=x^{2}+4y^{2},$
where $x$ and $y$ are integers, by [8, Eq. (51)]. Due to [8, Eq. (55)], we have
(5) $x=2q-1-8(1,0)_{4},$
where $(1,0)_{4}$ is a cyclotomic number modulo $4$ (see [8] for the definition).
If $|D|=q$, then $(1,0)_{4}=\frac{q-1}{4}$ by [3, VI.Theorem 8.10(a)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q+2=1$. Therefore $p=4q+1=4y^{2}+1$ by Eq. (4) and hence $q=y^{2}$, a contradiction to primality of $q$. If $|D|=q+1$, then $(1,0)_{4}=\frac{q+1}{4}$ by [3, VI.Theorem 8.10(b)]. Together with Eq. (5), this implies that $x=2q-1-8(1,0)_{4}=2q-1-2q-2=-3$. Eq. (4) yields that $p=4q+1=4y^{2}+9$ and hence $q=y^{2}+2$. If $y^{2}\equiv 1\mod~{}3$, then $q$ is divided by $3$. So $q=3$, a contradiction to the assumption of the theorem. If $y^{2}$ is divided by $3$, then $q\equiv 2\mod~{}3$. Therefore $p=4q+1$ is divided by $3$, a contradiction to primality of $p$. ∎
Remark 1.
If $q=3$, then there is a difference set in $A$ with parameters $(4q+1,q+1,\frac{q+1}{4})=(13,4,1)$ which is a Singer difference set (see [2]).
Let $G$ be a dihedral group, $A$ a cyclic subgroup of $G$ of index $2$, and $b$ an involution from $G\setminus A$. Suppose that $D$ is a difference set in $A$. Then [26, Lemma 5.1] implies that the partition of $G$ into sets
$\{e\},A^{\#},Db,(A\setminus D)b$
defines the $S$-ring $\mathcal{A}=\mathcal{A}(D)$. Clearly, $\operatorname{rk}(\mathcal{A}(D))=4$ and $\mathcal{A}(D)=\mathcal{A}(A\setminus D)$.
4. $S$-rings over $D_{2p}$
A cyclic and dihedral groups of order $n$ are denoted by $C_{n}$ and $D_{n}$, respectively. Throughout this section, $p$ is an odd prime, $G=\langle a,b:~{}a^{p}=b^{2}=e,~{}a^{b}=a^{-1}\rangle\cong D_{2p}$, $A=\langle a\rangle\cong C_{p}$, and $B=\langle b\rangle\cong C_{2}$. In this section, we classify $S$-rings over some dihedral groups. More precisely, we prove the following theorem.
Theorem 4.1.
Let $p$ be a prime. Suppose that $p$ is a Fermat prime or $p=rq+1$, where $q$ is a prime and $r\in\{2,4\}$. Then for every $S$-ring $\mathcal{A}$ over $D_{2p}$, one of the following statements holds:
  1. (1)
    $\operatorname{rk}(\mathcal{A})=2$;
  2. (2)
    $\mathcal{A}$ is cyclotomic;
  3. (3)
    $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$;
  4. (4)
    $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$;
  5. (5)
    $p\in\{13,2q+1\}$ and $\mathcal{A}=\mathcal{A}(D)$ for some nontrivial difference set $D$ in $A$.
It should be mentioned that every $S$-ring over $C_{2p}$ is of rank $2$, or cyclotomic, or a wreath product of two $S$-rings by [9, Theorem 4.1-4.2] (see also [20, 21]). The proof of Theorem 4.1 will be given in the end of the section. We say that a subset $X$ of $G$ is mixed if $X\cap A\neq\varnothing$ and $X\cap Ab\neq\varnothing$; otherwise we say that $X$ is nonmixed. Throughout this section, $\mathcal{A}$ is an $S$-ring over $G$. The next lemma immediately follows from the fact that $G$ is a strong B-group (see [34, Theorem 26.6(i)]).
Lemma 4.2.
If $\mathcal{A}$ is primitive, then $\operatorname{rk}(\mathcal{A})=2$.
Further we assume that $\mathcal{A}$ is imprimitive.
Lemma 4.3.
If $G$ has an $\mathcal{A}$-subgroup $L$ of order $2$, then $\mathcal{A}$ is cyclotomic or isomorphic to an $S$-ring over $C_{2p}$.
Proof.
Let $L=\{e,s\}$. Then $\{s\}\in\mathcal{S}(\mathcal{A})$. Suppose that $A$ is an $\mathcal{A}$-subgroup. Lemma 2.1 implies that $\mathcal{A}_{A}=\operatorname{Cyc}(K_{0},A)$ for some $K_{0}\leq\operatorname{Aut}(A)$. Every basic set of $\mathcal{A}$ outside $A\cup L$ is of the form $Xs$, where $X\in\mathcal{S}(\mathcal{A}_{A})=\operatorname{Orb}(K_{0},A)$. Therefore $\mathcal{A}=\operatorname{Cyc}(K,G)$, where $K\leq\operatorname{Aut}(G)$ is such that the restriction of $K$ on $A$ is equal to $K_{0}$ and $K$ fixes $s$. Further we assume that $A$ is not an $\mathcal{A}$-subgroup.
Assume that there exists a nonmixed basic set $Y$ of $\mathcal{A}$ outside $L$. If $Y\subseteq A$, then $A=\langle Y\rangle$ is an $\mathcal{A}$-subgroup, a contradiction to the assumption. If $Y\subseteq Ab$, then $Ys$ is a basic set inside $A$ and $A=\langle Ys\rangle$ is an $\mathcal{A}$-subgroup, a contradiction to the assumption.
Due to the above paragraph, we may assume that every basic set $Y$ of $\mathcal{A}$ outside $L$ is mixed and hence of the form $Y=Y_{0}\cup Y_{1}s$, where $Y_{0}$ and $Y_{1}$ are nonempty subsets of $A$. Note that $Y_{1}s=(Y_{1}s)^{-1}$. So $Y=Y^{-1}$ and $Y_{0}=Y_{0}^{-1}$. Since $\{s\}\in\mathcal{S}(\mathcal{A})$, the set $Ys=Y_{0}s\cup Y_{1}$ is basic. Moreover, $Ys=(Ys)^{-1}$ and hence $Y_{1}=Y_{1}^{-1}$. Therefore $\sigma\in\operatorname{Aut}(G)$ such that $\sigma:(a,s)\mapsto(a^{-1},s)$ is a Cayley automorphism of $\mathcal{A}$. One can verify straightforwardly that $|\sigma b_{r}|=2$ and $a_{r}\sigma b_{r}=\sigma b_{r}a_{r}$. This implies that the group $K=\langle a_{r},\sigma b_{r}\rangle$ is isomorphic to $C_{2p}$. It is easy to see that $K$ is transitive on $G$ and hence regular. Thus, $\operatorname{Aut}(\mathcal{A})$ has a regular subgroup isomorphic to $C_{2p}$. This yields that $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$. ∎
Further we assume that
[…]
In this case $Ab$ is an $\mathcal{A}$-set. Due to Lemma 2.1, we have $\mathcal{A}_{A}=\operatorname{Cyc}(K,A)$ for some $K\leq\operatorname{Aut}(A)\cong C_{p-1}$. Let $|K|=m$. Clearly, $m$ divides $p-1$ and every nontrivial basic set of $\mathcal{A}_{A}$ is of size $m$.
Lemma 4.4.
If $Y\in\mathcal{S}(\mathcal{A})_{Ab}$, then $|Y|\notin\{1,p-1\}$.
Proof.
Obviously, $|Y|\leq|Ab|=p$. If $|Y|=1$ ($|Y|=p-1$, respectively), then $\langle Y\rangle$ ($\langle Ab\setminus Y\rangle$, respectively) is an $\mathcal{A}$-subgroup of order $2$, a contradiction to the assumption that $A$ is the unique $\mathcal{A}$-subgroup. ∎
Lemma 4.5.
Let $r$ be a prime and $s$ a positive integer. Suppose that $r^{s}$ divides $m$. Then there exists a unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r^{s}$ does not divide $|Y|$. Moreover, $|Y|\equiv 1~{}\mod~{}r^{s}$.
Proof.
Let $Y_{1},Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$. Then $Y_{1}=Y_{1}^{-1}$ and $Y_{2}=Y_{2}^{-1}$. If $Y_{1}=Y_{2}$, then
$\underline{Y_{1}}\cdot\underline{Y_{1}}=|Y_{1}|e+\sum_{\begin{subarray}{c}X\in\mathcal{S}(\mathcal{A})_{A}\\ X\neq\{e\}\end{subarray}}c_{Y_{1}Y_{2}}^{X}\underline{X},$
and if $Y_{1}\neq Y_{2}$, then
$\underline{Y_{1}}\cdot\underline{Y_{2}}=\sum_{\begin{subarray}{c}X\in\mathcal{S}(\mathcal{A})_{A}\\ X\neq\{e\}\end{subarray}}c_{Y_{1}Y_{2}}^{X}\underline{X}.$
Counting the number of elements in the left-hand side and the right-hand side of the above equalities and taking into account that $|X|=m$ for every nontrivial $X\in\mathcal{S}(\mathcal{A}_{A})$, we obtain
(6) $m~{}|~{}|Y_{1}|(|Y_{1}|-1)~{}\text{and}~{}m~{}|~{}|Y_{1}||Y_{2}|.$
Observe that $\sum\limits_{Y\in\mathcal{S}(\mathcal{A})_{Ab}}|Y|=|Ab|=p$. This implies that there exists $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $r$ does not divide $|Y|$; for otherwise $r$ divides $p$ which is impossible. The second part of Eq. (6) yields that $|Y||Z|$ is divided by $r^{s}$ and hence $|Z|$ is divided by $r^{s}$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. From the first part of Eq. (6) it follows that $|Y|-1$ is divided by $r^{s}$ or, equivalently, $|Y|\equiv 1~{}\mod~{}r^{s}$. ∎
Lemma 4.6.
If $\operatorname{GCD}(|Y|,m)=1$ for some $Y\in\mathcal{S}(\mathcal{A})_{Ab}$, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$.
Proof.
From Lemma 4.5 it follows that $|Z|$ is divided by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ with $Z\neq Y$. Note that $|Y|\geq 2$ by Lemma 4.4. Let $y_{1},y_{2}\in Yb$ such that $y_{1}\neq y_{2}$ and $X$ a nontrivial basic set of $\mathcal{A}_{A}$ such that $y_{2}y_{1}^{-1}\in X$. Since $X\subseteq A$ and $Y\subseteq Ab$, we have
(7) $\underline{X}\cdot\underline{Y}=c_{XY}^{Y}\underline{Y}+\sum_{\begin{subarray}{c}Z\in\mathcal{S}(\mathcal{A})_{Ab},\\ Z\neq Y\end{subarray}}c_{XY}^{Z}\underline{Z}.$
Counting the number of elements in the left-hand side and the right-hand side of the above equality and taking into account that $|X|=m$ and $|Z|$ is divided by $m$ for every $Z\in\mathcal{S}(\mathcal{A})_{Ab}$ distinct from $Y$, we conclude that
(8) $m|Y|=c_{XY}^{Y}|Y|+\alpha m$
for some nonnegative integer $\alpha$. Since $|Y|$ and $m$ are coprime, the number $c_{XY}^{Y}$ is divided by $m$ due to Eq. (8). Observe that $c_{XY}^{Y}\geq 1$ because $y_{2}y_{1}^{-1}\in X$ and the element $y_{2}b$ enters the element $\underline{X}\cdot\underline{Y}$ with nonzero coefficient. So $c_{XY}^{Y}\geq m$. Now Eq. (8) yields that $c_{XY}^{Y}=m$ and $\alpha=0$. Therefore from Eq. (7) it follows that $\underline{X}\cdot\underline{Y}=|X|\underline{Y}$ and hence $X\leq\operatorname{rad}(Y)$. Thus, $A=\langle X\rangle\leq\operatorname{rad}(Y)$ and $Y=Ab$. The latter means that $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$. ∎
Lemma 4.7.
If $m$ is a prime power, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$.
Proof.
If $m=1$, then we are done by Lemma 4.6. If $m>1$, then there exists the unique $Y\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $\operatorname{GCD}(|Y|,m)=1$ by Lemma 4.5 because $m$ is a prime power. Again, we are done by Lemma 4.6. ∎
Lemma 4.8.
If $p=rq+1\geq 29$, where $q$ is a prime and $r\in\{2,4\}$, and $m\neq p-1$, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$.
Proof.
Note that $m$ divides $p-1$. If $m\in\{1,2,4,q\}$, then we are done by Lemma 4.7. Since $m\neq p-1$, it remains to consider the case $r=4$ and $m=2q$. In this case, $\operatorname{rk}(\mathcal{A}_{A})=3$. Let $X_{1}$ and $X_{2}$ be nontrivial basic sets of $\mathcal{A}_{A}$. Clearly, $|X_{1}|=|X_{2}|=2q$. Since $m$ is even, we have $X_{1}=X_{1}^{-1}$ and $X_{2}=X_{2}^{-1}$. Thus, $\mathcal{A}$ is symmetric and hence commutative.
Lemma 4.5 implies that there exist a unique $Y_{1}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{1}|$ is not divided by $q$ and a unique $Y_{2}\in\mathcal{S}(\mathcal{A})_{Ab}$ such that $|Y_{2}|$ is not divided by $2$. If $Y_{1}=Y_{2}$, then $\operatorname{GCD}(|Y_{1}|,m)=1$ and we are done by Lemma 4.6.
Suppose that $Y_{1}\neq Y_{2}$. From Lemma 4.5 it follows that $|Y_{1}|\equiv 1\mod q$. Since $Y_{2}$ is a unique basic set from $\mathcal{S}(\mathcal{A})_{Ab}$ of odd cardinality and $Y_{1}\neq Y_{2}$, we conclude that $|Y_{1}|$ is divided by $2$. Therefore
$|Y_{1}|\in\{q+1,3q+1\}.$
Let $|Y_{1}|=3q+1$. Since $|Y_{2}|$ is divided by $q$, we have $\mathcal{S}(\mathcal{A})_{Ab}=\{Y_{1},Y_{2}\}$ and $|Y_{2}|=q$ in this case. Note that
(9) $c_{Y_{1}Z}^{T}=0$
whenever $Z,T\in\mathcal{S}(\mathcal{A}_{A})$ or $Z,T\in\mathcal{S}(\mathcal{A})_{Ab}$. So by Eq. (2), we have
$|Y_{1}|=c_{Y_{1}\{e\}}^{Y_{1}}+c_{Y_{1}X_{1}}^{Y_{1}}+c_{Y_{1}X_{2}}^{Y_{1}}=1+c_{Y_{1}X_{1}}^{Y_{1}}+c_{Y_{1}X_{2}}^{Y_{1}}.$
Therefore
(10) $c_{Y_{1}X_{1}}^{Y_{1}}+c_{Y_{1}X_{2}}^{Y_{1}}=|Y_{1}|-1.$
Eq. (1) and symmetry of $\mathcal{A}$ imply that
$c_{Y_{1}X_{1}}^{Y_{1}}|Y_{1}|=c_{Y_{1}Y_{1}}^{X_{1}}|X_{1}|~{}\text{and}~{}c_{Y_{1}X_{2}}^{Y_{1}}|Y_{1}|=c_{Y_{1}Y_{1}}^{X_{2}}|X_{2}|.$
Since $|Y_{1}|$ is not divided by $q$ and $|X_{1}|=|X_{2}|=2q$, the above equalities yield that $c_{Y_{1}X_{1}}^{Y_{1}}$ and $c_{Y_{1}X_{2}}^{Y_{1}}$ are divided by $q$. If $|Y_{1}|=3q+1$, then $c_{Y_{1}X_{1}}^{Y_{1}}\geq 2q$ or $c_{Y_{1}X_{2}}^{Y_{1}}\geq 2q$ by Eq. (10). The latter means that $Y_{1}X_{1}=Y_{1}$ or $Y_{1}X_{2}=Y_{1}$. In both cases $\operatorname{rad}(Y_{1})$ is nontrivial and hence $Y_{1}=Ab$, a contradiction to $Y_{1}\neq Y_{2}$.
Now suppose that $|Y_{1}|=q+1$. Then Eq. (10) implies that $c_{Y_{1}X_{1}}^{Y_{1}}=0$ or $c_{Y_{1}X_{2}}^{Y_{1}}=0$. Without loss of generality, let $c_{Y_{1}X_{1}}^{Y_{1}}=c_{X_{1}Y_{1}}^{Y_{1}}=0$. Put $Z_{1}=Y_{1}b\subseteq A$. Since $c_{X_{1}Y_{1}}^{Y_{1}}=0$, there is no an element from $Z_{1}$ which enters the element $\underline{X_{1}}\cdot\underline{Z_{1}}$. Therefore
(11) $|Z_{1}\cap X_{1}z|=0$
for every $z\in Z_{1}$. The Cayley graph $\operatorname{Cay}(A,X_{1})$ is a Paley graph with parameters $(4q+1,2q,q-1,q)$. Due to Eq. (11), any two elements from $Z_{1}$ are non-adjacent in $\operatorname{Cay}(A,X_{1})$. This yields that
(12) $|X_{1}z\cap X_{1}z^{\prime}|=q$
for any $z,z^{\prime}\in Z_{1}$. Clearly, $|Z_{1}|=|Y_{1}|=q+1\geq 8$. Let $z,z^{\prime},z^{\prime\prime}\in Z_{1}$ be pairwise distinct. Since $p\geq 29$,
(13) $|X_{1}z\cap X_{1}z^{\prime}\cap X_{1}z^{\prime\prime}|\geq 1.$
by [4, Corollary 0.2]. Using Eqs. (11), (12), (13), we obtain
$|Z_{1}\cup X_{1}z\cup X_{1}z^{\prime}\cup X_{1}z^{\prime\prime}|=|Z_{1}|+|X_{1}z|+|X_{1}z^{\prime}|+|X_{1}z^{\prime\prime}|-|Z_{1}\cap X_{1}z|-|Z_{1}\cap X_{1}z^{\prime}|-|Z_{1}\cap X_{1}z^{\prime\prime}|-$
$-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z\cap X_{1}z^{\prime}|-|X_{1}z\cap X_{1}z^{\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime}|+|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime\prime}|+$
$+|Z_{1}\cap X_{1}z^{\prime}\cap X_{1}z^{\prime\prime}|+|X_{1}z\cap X_{1}z^{\prime}\cap X_{1}z^{\prime\prime}|-|Z_{1}\cap X_{1}z\cap X_{1}z^{\prime}\cap X_{1}z^{\prime\prime}|\geq$
$\geq q+1+6q-3q+1=4q+2>|A|,$
a contradiction to $Z_{1}\cup X_{1}z\cup X_{1}z^{\prime}\cup X_{1}z^{\prime\prime}\subseteq A$. ∎
Proof of Theorem 4.1.
The theorem for $p\in\{5,7,11,13,17,23\}$ follows from computer calculations (see [36]). Further, we assume that $p\geq 29$.
If $\mathcal{A}$ is primitive, then $\operatorname{rk}(\mathcal{A})=2$ by Lemma 4.2 and Statement $(1)$ of Theorem 4.1 holds. Suppose that $\mathcal{A}$ is imprimitive. If there is an $\mathcal{A}$-subgroup $L$ of order $2$, then $\mathcal{A}$ is cyclotomic or $\mathcal{A}$ is isomorphic to an $S$-ring over $C_{2p}$ by Lemma 4.3 and Statement $(2)$ or $(3)$ of Theorem 4.1 holds.
Now suppose that there is no an $\mathcal{A}$-subgroup of order $2$. Since $\mathcal{A}$ is imprimitive, $A$ is a unique $\mathcal{A}$-subgroup. If $p$ is a Fermat prime, then $m$ is a $2$-power and $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 4.7 and Statement $(4)$ of Theorem 4.1 holds. Let $p=rq+1$, where $q$ is a prime and $r\in\{2,4\}$. We may assume that $q$ is odd because otherwise $p$ is a Fermat prime and we are done. If $m\neq p-1$, then $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$ by Lemma 4.8 and Statement $(4)$ of Theorem 4.1 holds.
Let $m=p-1$. Then $\operatorname{rk}(\mathcal{A}_{A})=2$ and $X=A^{\#}$ is a basic set of $\mathcal{A}_{A}$. Let $Y\in\mathcal{S}(\mathcal{A})_{Ab}$. If $|Y|=p$, then $Y=Ab$ and $\mathcal{A}=\mathcal{A}_{A}\wr\mathcal{A}_{G/A}$. So in view of Lemma 4.4, we may assume that $22\leq|Y|\leq p-$ 2. Since $\mathcal{A}$ is an $S$-ring,
$\underline{Y}\cdot\underline{Y}=\underline{Yb}\cdot(\underline{Yb})^{-1}=|Y|e+\lambda\underline{A}^{\#}$
for some positive integer $\lambda$. Therefore $D=Yb$ is a nontrivial difference set in $A$. If $p=4q+1$, we obtain a contradiction to Theorem 1.2. Suppose that $p=2q+1$. Due to the above discussion and Lemma 3.1, every basic set of $\mathcal{A}$ outside $A$ has size $q$ or $q+1$. Therefore $\operatorname{rk}(\mathcal{A})=4$ and hence $\mathcal{A}=\mathcal{A}(D)$. Thus, Statement $(5)$ of Theorem 4.1 holds. ∎
Proof of Theorem 1.1.
Let $p$ be a Fermat prime or $p=4q+1$, where $q$ is prime. Let us prove that $\mathcal{A}$ is schurian. One of the statements of Theorem 4.1 holds for $\mathcal{A}$. If Statement $(1)$ or $(2)$ holds, then $\mathcal{A}$ is obviously schurian. If Statement $(3)$ holds, then $\mathcal{A}$ is schurian because $C_{2p}$ is a Schur group by [10, Theorem 1.1]. If Statement $(4)$ holds, then $\mathcal{A}$ is schurian by Lemma 2.1 and Lemma 2.2. If Statement $(5)$ holds, then $p=13$ and $\mathcal{A}$ is schurian by computer calculations [36]. ∎
5. Further discussion
We do not know whether it is possible to extend Theorem 1.1 to other primes $p$. Recall that if $p\equiv 3\mod~{}4$ and $p>11$, then $D_{2p}$ is not Schur by [26, Corollary 5.3]. In this case, there is a Paley difference $D$ set in $C_{p}$ and the $S$-ring $\mathcal{A}(D)$ is nonschurian whenever $p>11$. However, $D_{2p}$ can be not Schur even if $p\equiv 1\mod~{}4$ which follows from the lemma below.
Lemma 5.1.
If $p$ is a prime of the form $p=4t^{2}+1$ or $p=4t^{2}+9$, where $t\geq 3$ is an odd integer, then $D_{2p}$ is not Schur.
Proof.
From [3, Theorem VI.8.11(b)] it follows that there is a biquadric residue difference set $D$ in a cyclic group of order $p$ with parameters $(v,k,\lambda)=(4t^{2}+1,t^{2},\frac{t^{2}-1}{4})$ if $p=4t^{2}+1$ and $(v,k,\lambda)=(4t^{2}+9,t^{2}+3,\frac{t^{2}+3}{4})$ if $p=4t^{2}+9$. Suppose that $p=4t^{2}+1$. Then $k=t^{2}$ divides $v-1=4t^{2}$. If $t>3$, then $41+\sqrt{k}=1+t>\frac{v-1}{k}=$ 4. So the design corresponding to $D$ is not $2$-transitive by [18, Theorem 8.3]. Therefore the $S$-ring over $D_{2p}$ constructed from $D$ as in [26, Lemma 5.1] is nonschurian by [26, Lemma 5.2] and hence $D_{2p}$ is not Schur. If $t=3$, then it is easy to see that $v=4t^{2}+1=37$ can not be presented in the form $\frac{q^{d+1}-1}{q-1}$ for some prime power $q$ and $d\geq 2$ and hence $D_{2p}$ is not Schur by [26, Corollary 5.4].
Now let $p=4t^{2}+9$. Assume that $D_{2p}$ is Schur. Then [26, Corollary 5.4] implies that $D$ must have parameters
$\left(\frac{q^{d+1}-1}{q-1},\frac{q^{d}-1}{q-1},\frac{q^{d-1}-1}{q-1}\right),$
where $q$ is a prime power and $d\geq 2$. The straightforward computation yields that
$\frac{k}{\lambda}=4=q+\frac{q-1}{q^{d-1}-1}.$
Therefore $q=3$, $d=2$, and $t=1$, a contradiction to $t\geq 3$. ∎
Observe that if $p=4t^{2}+9$, then $p-1$ can be of the form $p-1=4m$, where $m$ is odd square-free. For example, $p=3373=4\cdot 29^{2}+9=4\cdot 3\cdot 281+1$. It would be interesting to study schurity of $D_{2p}$ for all primes $p$ such that $p\equiv 1\mod~{}4$.